OpenStudy (anonymous):

find out the binary substraction by using binary complemetry method 10011101 from 11101000

OpenStudy (anonymous):

If I understand correctly, you are looking for 11101000-10011101 using the complementary method. Well, 11101000 +01100010 (since 10011101+01100010=11111111) + 1 _________ 101001011 now, subtract 100000000 making your answer 01001011 Basically what you are doing is adding a specific number (the radix I believe) to make the subtraction process easier. To check our answer, we can convert back to decimal so: 11101000 = 232 -01100010 = -157 ______________ 01001011 = 75 If it helps, this is how the process looks when you convert all the numbers to decimal first and then use the complementary method: 232 + 98 + 1 ______ 331 - 256 _____ 75